# Towards Robust Surrogate Models: Benchmarking Machine Learning Approaches to Expediting Phase Field Simulations of Brittle Fracture

### v1

@@ Abstract
Data driven approaches have the potential to make modeling complex, nonlinear physical phenomena significantly more computationally tractable. For example, computational modeling of fracture is a core challenge where machine learning techniques have the potential to provide a much needed speedup that would enable progress in areas such as mutli-scale modeling and uncertainty quantification. Currently, phase field modeling (PFM) of fracture is one such approach that offers a convenient variational formulation to model crack nucleation, branching and propagation. To date, machine learning techniques have shown promise in approximating PFM simulations. However, most studies rely on overly simple benchmarks that do not reflect the true complexity of the fracture processes where PFM excels as a method. To address this gap, we introduce a challenging dataset based on PFM simulations designed to benchmark and advance ML methods for fracture modeling. This dataset includes three energy decomposition methods, two boundary conditions, and 1,000 random initial crack configurations for a total of 6,000 simulations. Each sample contains 100 time steps capturing the temporal evolution of the crack field. Alongside this dataset, we also implement and evaluate Physics Informed Neural Networks (PINN), Fourier Neural Operators (FNO) and UNet models as baselines, and explore the impact of ensembling strategies on prediction accuracy. With this combination of our dataset and baseline models drawn from the literature we aim to provide a standardized and challenging benchmark for evaluating machine learning approaches to solid mechanics. Our results highlight both the promise and limitations of popular current models, and demonstrate the utility of this dataset as a testbed for advancing machine learning in fracture mechanics research.
@@ 1 Introduction
In recent years, data-driven modeling has gained significant traction across science and engineering due to its ability to approximate complex physical systems at a fraction of the computational cost of traditional high fidelity simulations [1, 2] for autonomous and non-autonomous systems where the input space varies with time [3]. By extracting patterns from data and learning compact representations, neural network based approximations can offer fast and flexible alternatives to conventional solvers. A wide range of architectures have been proposed to introduce inductive biases relevant to the underlying structure of physical problems. For instance, convolutional neural networks (CNNs) have been applied to structured grids [4] while graph neural networks (GNNs) have been applied to problems defined on unstructured meshes and irregular underlying structures [5, 6]. To better integrate governing physics into the learning process, physics informed neural networks (PINNs) [7] were introduced. These models use neural networks to parameterize the solution of partial differential equations (PDEs) and incorporate the residuals of the governing equations directly into the loss function. PINNs have been applied to a broad range of problems [8, 9, 10, 11, 12], but face scalability challenges as they typically need to be retrained for every change in initial or boundary condition. To overcome these limitations, operator learning approaches have emerged. These methods aim to learn mappings between infinite dimensional input and output function spaces, enabling generalization across different input space parameters such as initial and boundary conditions. Both purely data driven and physics constrained variants have been proposed within this framework achieving state of the art performance across a wide range of scientific and engineering domains [13, 14, 15, 16, 17, 18].
@@
However, despite these advances, neural network based solutions continue to face major challenges. For example, the models mentioned previously, apply the governing physical laws as soft penalties within the loss function rather than enforcing them strictly. As a result, the learned solutions may violate the underlying physics especially in regions with sparse data or high complexity [19, 20, 21, 22]. Additionally, many models suffer from what is known as spectral bias which is the tendency of the model to learn low frequency (smooth) components of the solution first while struggling to capture high frequency features [23, 24]. This makes it challenging for these models to accurately capture regions with sharp gradients, such as discontinuities or localized damage. This behavior is particularly problematic for modeling fracture where small scale features such as cracks play a critical role in determining the system’s overall response. Other persistent issues include difficulty of optimizing the neural networks in a non convex loss landscape [25, 26, 27] and sensitivity to initialization or training dynamics [28, 29, 30]. Together, these limitations make it difficult to develop robust, generalizable and physically faithful models highlighting the need for structured, high fidelity benchmark datasets to systematically study and stress test model capabilities such as the Mechanical MNIST collection [31, 32, 33] which presented accessible benchmark datasets for mechanical engineering inspired from the computer vision research community or other benchmarks that target broader scientific meta modeling benchmarks such as the Well [34] or PDEBench [35].
@@
Inspired by these challenges, we have chosen brittle fracture as the starting point for our dataset. There are many different ways to model brittle fracture [36, 37, 38]. Among these methods, Phase field modeling (PFM) of brittle fracture is a popular modern option that has been explored recently within the ML literature [39, 40]. Notably fracture problems are inherently characterized by discontinuities, sharp gradients, and topological changes such as crack nucleation, branching, and coalescence, which, as mentioned, are particularly difficult for data driven models to capture accurately. At the same time, the coupled governing equations of the PFM provides a well defined and physically grounded framework with intrinsic non convexity [41] making it a natural setting to embed physics into neural networks and rigorously evaluate their robustness and convergence behavior under complex physical settings. PFMs have also been explored in the context of stochastic modeling [41] and uncertainty quantification [42], and are notorious for being computationally expensive [43, 44, 45]. Different attempts have been made by researchers to offer a data driven approximation to the solution of phase field models of brittle fracture [46, 47, 40, 48, 17]. However many of these studies have ultimately evaluated ML models on relatively simple benchmark problems. Without a rigorous exploration of more complex and diverse fracture scenarios, it remains unclear whether these ML models are truly able to capture the subtleties of fracture prediction.
@@
In this work our primary goal is to develop a challenging dataset that enables a deeper exploration of the subtleties of surrogate modeling with a particular focus on phase field modeling of brittle fracture. As can be seen in Fig. 1 this dataset is designed not only to reflect the complex nature of fracture and crack propagation under various boundary conditions and energy decompositions, but also to serve as a robust benchmark for evaluating the performance of data driven surrogate models. Specifically, we implement and assess the performance of several baseline neural network based models including Physics informed neural networks (PINNs), UNets, and Fourier Neural Operators (FNOs) in order to better understand their strengths and limitations in modeling fracture. A further objective of this work is to evaluate the impact of three ensembling strategies (hard voting, soft voting and stacking) on the performance of these models, and to assess their effectiveness in improving predictive accuracy and robustness. This dataset extends our previous contributions to benchmark dataset development in mechanics, including the Mechanical MNIST dataset collection [31]. Specifically, it builds upon the Mechanical MNIST Crack Path dataset [49] by introducing a more advanced and more broadly relevant dataset. We refer to this new dataset, "PFM-Fracture" subset of our new "Nonlinear Mechanics Benchmark" dataset which consists of high-resolution finite element simulations of brittle fracture using PFM applied to domains with randomly initialized cracks.
@@
Figure 1: Comparison between approaches to evaluating surrogate model efficacy: a) Standard examples from the literature such as tension, shear, and crack coalescences [46, 40], b) Our proposed benchmark dataset designed to provide a diverse set of challenges across initial conditions, load conditions, and energy decomposition methods.
@@ 2.1.1 Phase Field Modeling of Fracture
$\mathcal{E}(\mathbf{u},\phi)=\Psi_{s}+\Psi_{c}-\mathcal{P}=\int_{\Omega}\psi_{0}(\mathbf{\mathbf{\epsilon}}(\mathbf{u}))\mathrm{d}V+\int_{\Gamma}G_{c}\mathrm{d}A-\int_{\Omega}\mathbf{b}\cdot\mathbf{u}\mathrm{d}V+\int_{\partial\Omega}\mathbf{t}\cdot\mathbf{u}\mathrm{d}A$ (1)
@@
$\Psi_{s}(u,\phi)=\int_{\Omega}g(\phi)\psi_{0}(\mathbf{\epsilon}(u))\mathrm{d}V$ (6)
@@
Various variational and non-variational approaches have been proposed in the literature. In variational formulations, the stress field and phase-field evolution equations can be derived from Eq.7, as seen in [56, 44]. However, these formulations lead to nonlinear equilibrium equations, which are computationally expensive to solve. To address this issue, hybrid formulations have been introduced [45, 57], incorporating a linear displacement subproblem to reduce the computational cost. Given the complexity of each simulation sample—particularly due to the need to resolve crack coalescence and the irregularity of initial conditions—we have implemented the hybrid phase field formulation proposed by Ambati et al. [45], which maintains a linear momentum equation to reduce computational costs. With this approach, the linear momentum equation is written as follows:
[…]
where the history field variable $\mathcal{H}^{+}$ is adopted from [44] to enforce the irreversibility of the crack phase field evolution where $\psi_{0}^{+}$ is the tensile component of elastic energy density. Each of the displacement field and phase-field functional are minimized iteratively in an alternating minimization framework, which offers a robust but slow convergence [43]. An important feature of Eq.7 is that it is convex with respect to $\mathbf{u}$ and $\phi$, but is not convex with respect to the both [41]. This further necessitates the use of an alternating minimization method for solving the PDEs [43]. To enforce tensile-compressive asymmetry in the fracture response, the elastic energy density function must be decomposed into tensile and compressive parts where only the positive component should be degraded using the energy degradation function $g(\phi)=(1-\phi)^{2}$ as in Eq.9. With this approach, $\psi_{0}(\mathbf{\epsilon}(\mathbf{u}))$ is written as:
@@ 2.1.2 Energy Decomposition Methods
[…]
@@
Figure 3: Geometrical description and boundary conditions of the three benchmark cases and the resulting crack path and force displacement curves to validate the implemented phase-field model. Here we show: a) the tension case, b) the shear case and c) the crack coalescence case under uniaxial tension. The crack patterns for the spectral decomposition is depicted in (ii). The force displacement curve for all three of our energy decomposition approaches are reported in (iii) alongside results from the literature.
@@ 2.1.3 Benchmark Validation Problems
However, despite the fact that comparing to standard benchmarks is common practice, in our analysis of the literature we found that it was often quite challenging to reproduce the benchmark results reported by others. Critically, phase field fracture approaches are sensitive to multiple modeling parameters, including basic mesh structure, characteristic length, convergence criteria, and the strategy for modeling pre-exsiting cracks [41, 43]. And, because sharing software and implementation code are not standard practice, it is especially difficult to isolate sources of discrepancy. To this end, we make a note of our implementation choices in Section 2.2, and share all of the code underlying our implementation of the models shown in this Section. To validate our implementation of the three approaches outlined in Section 2.1.2 under different types of loading and boundary conditions, we compare their performance to each other and to the literature on three standard benchmark problems. These popular benchmark cases include: (1) Tension loading on a single-edged notched specimen – referred to as the “tension case,” (2) shear loading on a single-edged notched specimen – referred to as the “shear case,” and (3) uniaxial tension loading on a domain with internal flaws where crack coalescence is expected – referred to as the “crack coalescence case.” We show our results in Fig. 3 and provide additional details below. As briefly mentioned in Section 2.1.1, an extremely fine mesh is required in the crack transition zones to accurately compute the crack surface energy. The literature suggests that the mesh size in the cracked region should be at least $h=l_{0}/2$ as suggested in [70, 44]. Another important aspect in implementing PFM models is the definition of initial pre-existing cracks. Consistent with previous reports from the literature [43], different methods for defining the initial crack can result in a different force displacement response, which may explain small discrepancies between our force-displacement curve results and the results from other implementations. In this work we have used the initial strain history defined in [70]
@@
Here, $\phi_{c}=0.999$ represents the phase-field scalar value assigned to pre-existing cracks, while $d(x)$ denotes the shortest distance between points within the domain and the crack.
@@
The tension case domain is illustrated in Fig. 3 a-i. Specifically, it is a $1\times 1$ mm square containing an initial crack with length 0.5 mm in the middle of the left edge. The top boundary experiences a uniform vertical displacement with a rate of $\Delta u=1E-5$ mm for the first 500 steps, and $\Delta u=1E-6$ mm for the next 1500 steps. For this case, we use an unstructured mesh with linear triangular elements refined in the central band of the domain where the crack is expected to propagate. The smallest element size inside the region is $h=0.001$ mm, and the largest element size outside the crack zone is $h=0.01$ mm. The convergence criteria tolerance value was set to $41E-$ 4. The initial crack was modeled by using an initial strain history field following Borden et al. [70]. All other parameters were set to match the recent literature [45] with Young’s modulus set to $E=210$ GPa, $\nu=0.3$, $G_{c}=2.7$ N/mm and characteristic length $l_{0}=4E-3$ mm. As shown in Fig. 3 a-ii, where the typical resulting crack pattern is illustrated, and in Fig. 3 a-iii where the force displacement curves are illustrated, the results obtained from our simulations for the different energy decompositions are in good agreement with the literature [45].
@@
For the shear case, all physical parameters (Young’s modulus $E$, critical energy release rate $G_{c}$, characteristic length $l_{0}$ and the poisson’s ratio $\nu$) match the tension case. We use a uniform structured mesh with $256\times 256$ Q1 elements, following the original implementation by [45]. The top boundary is subject to shear with a constant rate of $\Delta u=1E-5$ mm. The boundary conditions are shown in Fig. 3 b-i , a representative resulting crack pattern is shown in Fig. 3 b-ii, and the load displacement curves are shown in Fig. 3 b-iii. Consistent with the literature [45], there are differences in the force displacement curve response using volumetric-deviatoric decomposition when compared to the other models. The spectral decomposition and the star-convex decomposition approaches show a similar response.
@@
Beyond the standard tension and shear benchmark cases, we also implemented a benchmark case from Zhou et al. [73] due to its similarity to the final dataset proposed in this work. In Zhou et al., initial flaws are modeled as geometry-induced cracks inside the mesh, but in our model we use a strain history field for defining the initial cracks. To closely match the force-displacement curve presented in Zhou et al. [73], we used a characteristic length of $l_{0}=0.12$ mm and a regionally refined unstructured mesh with triangular elements where the smallest element size within the region is $h=0.05$ mm and the largest mesh element size is $h=0.25$mm. As illustrated in Fig. 3 c-i, the domain is under uniaxial tension. Young’s modulus is set to $E=30$ GPa, $\nu=0.333$ and $G_{c}=3e-3$ N/mm. As can be seen in Fig. 3 c-ii-iii, the crack pattern and the force displacement curves are in good agreement with Zhou et al. [73] which uses the spectral decomposition approach. However, it is worth pointing out that in this case, using either the volumetric-deviatoric or the star-convex decomposition results in branching of the crack pattern near the boundaries as shown in the Fig. 4.
@@
Figure 6: Figure showing the initial crack pattern and final crack pattern for three samples of each dataset subcategory. From these illustrative results we see that in some cases almost no difference between the three energy decomposition methods (see $3^{rd}$ row of bi-axial tension case and $2^{nd}$ row of shear case), moderate difference ($1^{st}$ row of bi-axial tension case and shear case) or extensive difference ($3^{rd}$ row of shear case) in the resulting final crack pattern.

### v2

@@ Abstract
Data-driven approaches have the potential to make modeling complex, nonlinear physical phenomena significantly more computationally tractable. For example, computational modeling of fracture is a core challenge where machine learning techniques have the potential to provide a much needed speedup that would enable progress in areas such as multi-scale modeling and uncertainty quantification. Currently, phase field modeling (PFM) of fracture is one such approach that offers a convenient variational formulation to model crack nucleation, branching and propagation. To date, machine learning techniques have shown promise in approximating PFM simulations. While standard fracture benchmarks represent realistic scenarios frequently observed in practice, they typically do not provide sufficiently challenging tests for data-driven methods. To address this gap, we introduce a challenging dataset based on PFM simulations designed to benchmark and advance ML methods for fracture modeling. This dataset includes three energy decomposition methods, two boundary conditions, and 1,000 random initial crack configurations for a total of 6,000 simulations. Each sample contains 100 time steps capturing the temporal evolution of the crack field. Alongside this dataset, we also implement and evaluate Physics Informed Neural Networks (PINN), Fourier Neural Operators (FNO) and UNet models as baselines, and explore the impact of ensembling strategies on prediction accuracy. With this combination of our dataset and baseline models drawn from the literature we aim to provide a standardized and challenging benchmark for evaluating machine learning approaches to solid mechanics. Our results highlight both the promise and limitations of popular current models, and demonstrate the utility of this dataset as a testbed for advancing machine learning in fracture mechanics research.
@@ 1 Introduction
In recent years, data-driven modeling has gained significant traction across science and engineering due to its ability to approximate complex physical systems at a fraction of the computational cost of traditional high fidelity simulations [1, 2] for autonomous and non-autonomous systems where the input space varies with time [3]. This advantage is critical for challenges like uncertainty quantification, particularly in energy storage devices such as redox flow batteries [4]. In these systems, complex phenomena like fracture and crack propagation contribute to capacity fade and reduced lifespan [5]. The design process for these complex systems involves studying coupled phenomena across multiple scales, making the use of traditional computational methods prohibitively expensive. Data-driven approaches, therefore, offer a promising solution for achieving the speedups required for effective simulation [6]. By extracting patterns from data and learning compact representations, neural network based approximations can offer fast and flexible alternatives to conventional solvers. A wide range of architectures have been proposed to introduce inductive biases relevant to the underlying structure of physical problems. For instance, convolutional neural networks (CNNs) have been applied to structured grids [7] while graph neural networks (GNNs) have been applied to problems defined on unstructured meshes and irregular underlying structures [8, 9]. To better integrate governing physics into the learning process, physics informed neural networks (PINNs) [10] were introduced. These models use neural networks to parameterize the solution of partial differential equations (PDEs) and incorporate the residuals of the governing equations directly into the loss function. PINNs have been applied to a broad range of problems [11, 12, 13, 14, 15], but face scalability challenges as they typically need to be retrained for every change in initial or boundary condition. To overcome these limitations, operator learning approaches have emerged. These methods aim to learn mappings between infinite dimensional input and output function spaces, enabling generalization across different input space parameters such as initial and boundary conditions. Both purely data-driven and physics constrained variants have been proposed within this framework achieving state of the art performance across a wide range of scientific and engineering domains [16, 17, 18, 19, 20, 21].
@@
However, despite these advances, neural network based solutions continue to face major challenges. For example, the models mentioned previously, apply the governing physical laws as soft penalties within the loss function rather than enforcing them strictly. As a result, the learned solutions may violate the underlying physics especially in regions with sparse data or high complexity [22, 23, 24, 25]. Additionally, many models suffer from a phenomenon known as spectral bias which is the tendency of the model to learn low frequency (smooth) components of the solution first while struggling to capture high frequency features [26, 27]. This makes it challenging for these models to accurately capture regions with sharp gradients, such as discontinuities or localized damage. This behavior is particularly problematic for modeling fracture where small scale features such as cracks play a critical role in determining the system’s overall response. Other persistent issues include difficulty of optimizing the neural networks in a non-convex loss landscape [28, 29, 30] and sensitivity to initialization or training dynamics [31, 32, 33]. Together, these limitations make it difficult to develop robust, generalizable and physically faithful models highlighting the need for structured, high fidelity benchmark datasets to systematically study and stress test model capabilities such as the Mechanical MNIST collection [34, 35, 36] which presented accessible benchmark datasets for mechanical engineering inspired from the computer vision research community or other benchmarks that target broader scientific meta modeling benchmarks such as the Well [37] or PDEBench [38].
@@
Inspired by these challenges, we have chosen fracture as the starting point for our dataset. Fracture and crack propagation is a particularly critical failure mechanism in real engineering systems. For example, in energy storage devices such as lithium-ion batteries, imperfections such as pre-existing microcracks can propagate through an electrode and limit functionality. [5, 39, 40]. There are many different ways to model brittle fracture [41, 42, 43]. Among these methods, Phase field modeling (PFM) of brittle fracture is a popular modern option that has been explored recently within the ML literature [44, 45]. Notably fracture problems are inherently characterized by discontinuities, sharp gradients, and topological changes such as crack nucleation, branching, and coalescence, which, as mentioned, are particularly difficult for data-driven models to capture accurately. At the same time, the coupled governing equations of the PFM provide a well defined and physically grounded framework with intrinsic non-convexity [46] making it a natural setting to embed physics into neural networks and rigorously evaluate their robustness and convergence behavior under complex physical settings. PFMs have also been explored in the context of stochastic modeling [46] and uncertainty quantification [47], and are notorious for being computationally expensive [48, 49, 50]. Different attempts have been made by researchers to offer a data-driven approximation to the solution of phase field models of brittle fracture [51, 52, 45, 53, 20]. However many of these studies have ultimately evaluated ML models on relatively simple benchmark problems. Without a rigorous exploration of more complex and diverse fracture scenarios, it remains unclear whether these ML models are truly able to capture the subtleties of fracture prediction.
@@
In this work our primary goal is to develop a challenging dataset that enables a deeper exploration of the subtleties of surrogate modeling with a particular focus on phase field modeling of brittle fracture. As can be seen in Fig. 1 this dataset is designed not only to reflect the complex nature of fracture and crack propagation under various boundary conditions and energy decompositions, but also to serve as a robust benchmark for evaluating the performance of data-driven surrogate models. Specifically, we implement and assess the performance of several baseline neural network based models including Physics informed neural networks (PINNs), UNets, and Fourier Neural Operators (FNOs) in order to better understand their strengths and limitations in modeling fracture. A further objective of this work is to evaluate the impact of three ensembling strategies (hard voting, soft voting and stacking) on the performance of these models, and to assess their effectiveness in improving predictive accuracy and robustness. This is motivated by the fact that a significant challenge in training surrogate models on a finite number of data points is the introduction of epistemic uncertainty due to approximation errors [54]. While many approaches exist for quantifying uncertainty in neural networks [54, 55], ensembling methods based on random initialization offer a simple yet effective strategy. By exploring different regions of the non-convex loss landscape, ensembles can improve the predictive performance of the neural networks and provide a practical method for uncertainty quantification [56, 57, 58]. This dataset extends our previous contributions to benchmark dataset development in mechanics, including the Mechanical MNIST dataset collection [34]. Specifically, it builds upon the Mechanical MNIST Crack Path dataset [59] by introducing a more advanced and more broadly relevant dataset. We refer to this new dataset, "PFM-Fracture" subset of our new "Nonlinear Mechanics Benchmark" dataset which consists of high-resolution finite element simulations of brittle fracture using PFM applied to domains with randomly initialized cracks.
@@
Figure 1: Comparison between approaches to evaluating surrogate model efficacy: a) Standard examples from the literature such as tension, shear, and crack coalescence [51, 45], b) Our proposed benchmark dataset designed to provide a diverse set of challenges across initial conditions, load conditions, and energy decomposition methods.
@@ 2.1.1 Phase Field Modeling of Fracture
$\mathcal{E}(\mathbf{u},\phi)=\Psi_{s}+\Psi_{c}-\mathcal{P}=\int_{\Omega}\psi_{0}(\mathbf{\epsilon}(\mathbf{u}))\mathrm{d}V+\int_{\Gamma}G_{c}\mathrm{d}A-\int_{\Omega}\mathbf{b}\cdot\mathbf{u}\mathrm{d}V+\int_{\partial\Omega}\mathbf{t}\cdot\mathbf{u}\mathrm{d}A$ (1)
@@
$\Psi_{s}(\mathbf{u},\phi)=\int_{\Omega}g(\phi)\psi_{0}(\mathbf{\epsilon}(\mathbf{u}))\mathrm{d}V$ (6)
@@
Various variational and non-variational approaches have been proposed in the literature. In variational formulations, the stress field and phase field evolution equations can be derived from Eq.7, as seen in [66, 49]. However, these formulations lead to nonlinear equilibrium equations, which are computationally expensive to solve. To address this issue, hybrid formulations have been introduced [50, 67], incorporating a linear displacement subproblem to reduce the computational cost. Given the complexity of each simulation sample—particularly due to the need to resolve crack coalescence and the irregularity of initial conditions—we have implemented the hybrid phase field formulation proposed by Ambati et al. [50], which maintains a linear momentum equation to reduce computational costs. With this approach, the linear momentum equation is written as follows:
$\left\{\begin{aligned} \sigma(\mathbf{u},\phi):=&(1-\phi)^{2}\frac{\partial\psi_{0}(\mathbf{\epsilon})}{\partial\mathbf{\epsilon}},\\ -l_{0}^{2}\Delta\phi+\phi=&\frac{2l_{0}}{G_{c}}(1-\phi)\mathcal{H}^{+},\\ \mathcal{H}^{+}(x,t):=&\max_{\tau\in[0,t]}\psi_{0}^{+}\big(\mathbf{\epsilon}(x,\tau)\big)\end{aligned}\right$. (8)
where the history field variable $\mathcal{H}^{+}$ is adopted from [49] to enforce the irreversibility of the crack phase field evolution where $\psi_{0}^{+}$ is the tensile component of elastic energy density. Each of the displacement field and phase field functional are minimized iteratively in an alternating minimization framework, which offers a robust but slow convergence [48]. An important feature of Eq.7 is that it is convex with respect to $\mathbf{u}$ and $\phi$, but is not convex with respect to both [46]. This further necessitates the use of an alternating minimization method for solving the PDEs [48]. To enforce tensile-compressive asymmetry in the fracture response, the elastic energy density function must be decomposed into tensile and compressive parts where only the positive component should be degraded using the energy degradation function $g(\phi)=(1-\phi)^{2}$ as in Eq.9. With this approach, $\psi_{0}(\mathbf{\epsilon}(\mathbf{u}))$ is written as:
@@ 2.1.2 Energy Decomposition Methods
where the $\langle\rangle$ is defined as $\langle a\rangle:=\max\{a,0\}$, $\kappa_{0}=\lambda+2\mu/3$ denotes the bulk modulus, and $\mathbf{1}$ is the identity tensor. This approach offers a straightforward decomposition method that effectively captures key requirements for defining a new energy decomposition in the phase field method, as outlined in [68], while ensuring a feasible computational cost.
@@
Figure 3: Geometrical description and boundary conditions of the three benchmark cases and the resulting crack path and force displacement curves to validate the implemented phase field model. Here we show: a) the Tension case, b) the Shear case and c) the crack coalescence case under uniaxial tension. The crack patterns for the spectral decomposition is depicted in (ii). The force displacement curve for all three of our energy decomposition approaches are reported in (iii) alongside results from the literature.
@@ 2.1.3 Benchmark Validation Problems
However, despite the fact that comparing to standard benchmarks is common practice, in our analysis of the literature we found that it was often quite challenging to reproduce the benchmark results reported by others. Critically, phase field fracture approaches are sensitive to multiple modeling parameters, including basic mesh structure, characteristic length, convergence criteria, and the strategy for modeling pre-existing cracks [46, 48]. And, because sharing software and implementation code are not standard practice, it is especially difficult to isolate sources of discrepancy. To this end, we make a note of our implementation choices in Section 2.2, and share all of the code underlying our implementation of the models shown in this Section. To validate our implementation of the three approaches outlined in Section 2.1.2 under different types of loading and boundary conditions, we compare their performance to each other and to the literature on three standard benchmark problems. These popular benchmark cases include: (1) Tension loading on a single-edged notched specimen – referred to as the “Tension case”; (2) shear loading on a single-edged notched specimen – referred to as the “Shear case”; and (3) uniaxial tension loading on a domain with internal flaws where crack coalescence is expected – referred to as the “Crack coalescence case”. We show our results in Fig. 3 and provide additional details below. As briefly mentioned in Section 2.1.1, an extremely fine mesh is required in the crack transition zones to accurately compute the crack surface energy. The literature suggests that the mesh size in the cracked region should be at least $h\leq l_{0}/2$ as suggested in [80, 49]. Another important aspect in implementing PFM models is the definition of initial pre-existing cracks. Consistent with previous reports from the literature [48], different methods for defining the initial crack can result in a different force displacement response, which may explain small discrepancies between our force-displacement curve results and the results from other implementations. In this work we have used the initial strain history defined in [80]
@@
Here, $\phi_{c}=0.999$ represents the phase field scalar value assigned to pre-existing cracks, while $d(x)$ denotes the shortest distance between points within the domain and the crack.
@@
The tension case domain is illustrated in Fig. 3 a-i. Specifically, it is a $1\times 1$ mm square containing an initial crack with length 0.5 mm in the middle of the left edge. The top boundary experiences a uniform vertical displacement with a rate of $\Delta u=1\times 10^{-5}$ mm for the first 500 steps, and $\Delta u=1\times 10^{-6}$ mm for the next 1500 steps. For this case, we use an unstructured mesh with linear triangular elements refined in the central band of the domain where the crack is expected to propagate. The smallest element size inside the region is $h=0.001$ mm, and the largest element size outside the crack zone is $h=0.01$ mm. The convergence criteria tolerance value was set to $1\times 10^{-4}$. The initial crack was modeled by using an initial strain history field following Borden et al. [80]. All other parameters were set to match the recent literature [50] with Young’s modulus set to $E=210$ GPa, $\nu=0.3$, $G_{c}=2.7$ N/mm and characteristic length $l_{0}=4\times 10^{-3}$ mm. As shown in Fig. 3 a-ii, where the typical resulting crack pattern is illustrated, and in Fig. 3 a-iii where the force displacement curves are illustrated, the results obtained from our simulations for the different energy decompositions are in good agreement with the literature [50].
@@
For the Shear case, all physical parameters (Young’s modulus $E$, critical energy release rate $G_{c}$, characteristic length $l_{0}$ and the poisson’s ratio $\nu$) match the tension case. We use a uniform structured mesh with $256\times 256$ Q1 elements, following the original implementation by [50]. The top boundary is subject to shear with a constant rate of $\Delta u=1\times 10^{-5}$ mm. The boundary conditions are shown in Fig. 3 b-i , a representative resulting crack pattern is shown in Fig. 3 b-ii, and the load displacement curves are shown in Fig. 3 b-iii. Consistent with the literature [50], there are differences in the force displacement curve response using volumetric-deviatoric decomposition when compared to the other models. The spectral decomposition and the star-convex decomposition approaches show a similar response.
@@
Beyond the standard Tension and Shear benchmark cases, we also implemented a benchmark case from Zhou et al. [83] due to its similarity to the final dataset proposed in this work. In Zhou et al., initial flaws are modeled as geometry-induced cracks inside the mesh, but in our model we use a strain history field for defining the initial cracks. To closely match the force-displacement curve presented in Zhou et al. [83], we used a characteristic length of $l_{0}=0.12$ mm and a regionally refined unstructured mesh with triangular elements where the smallest element size within the region is $h=0.05$ mm and the largest mesh element size is $h=0.25$ mm. As illustrated in Fig. 3 c-i, the domain is under uniaxial tension where we apply uniform vertical displacement with a rate of $\Delta u=5\times 10^{-7}$ mm on the top and bottom boundary and fix the x component of the displacement on the bottom boundary. Young’s modulus is set to $E=30$ GPa, $\nu=0.333$ and $G_{c}=3\times 10^{-3}$ N/mm. As can be seen in Fig. 3 c-ii-iii, the crack pattern and the force displacement curves are in good agreement with Zhou et al. [83] which uses the spectral decomposition approach. However, it is worth pointing out that in this case, using either the volumetric-deviatoric or the star-convex decomposition results in branching of the crack pattern near the boundaries as shown in the Fig. 4. As a brief note, we ensured that we observed the same branching behavior for these two decompositions using a finer mesh than the one shown in Fig. 4 with $h=0.02$ mm and a smaller loading increment of $\Delta u=1\times 10^{-7}$ $m$.
@@
Figure 6: Figure showing the initial crack pattern and final crack pattern for three samples of each dataset subcategory. From these illustrative results we see that in some cases almost no difference between the three energy decomposition methods (see $3^{rd}$ row of the Bi-axial tension case and $2^{nd}$ row of the Shear case), moderate difference ($1^{st}$ row of Bi-axial tension case and Shear case) or extensive difference ($3^{rd}$ row of Shear case) in the resulting final crack pattern.